Probability questions
692 published questions with worked solutions.
- Given that x [0,1] and y [0,1]. Let A be the event of (x,y) satisfying y^2 x and B be the event of (x,y) satisfying x^2 y. Then
- The probability that exactly one of the independent events A and B occurs is equal to
- If A and B are independent events such that 0<P(A)<1, 0<P(B)<1 then
- For any two events A and B
- Four cards are drawn at random from a pack of 52 playing cards. The probability that the draw contains exactly one pair (i.e.,...
- If A and B are two independent events then which of the following is not equal to any of the remaining?
- An event X can take place in conjunction with any one of the mutually exclusive and exhaustive events A, B and C. If A, B, C are...
- Let E be an event which is neither a certainty nor an impossibility. Its probability is such that P(E)=1+ + ^ 2 and P(E')=(1+ )^...
- A number is selected at random from the first twenty-five natural numbers. If it is a composite number then it is divided by 5....
- A dice is thrown six times, it being known that each time a different digit shows up. The probability that a sum of 12 will be...
- A composite number is selected at random from the first 30 natural numbers and it is divided by 5. The probability that there...
- Two numbers are selected one by one without replacement at random from the set S= 1,2,3,4,5,6 . The probability that the smaller...
- Events A,B,C are mutually exclusive events such that P(A)= 3x+1 3 , P(B)= 1-x 4 and P(C)= 1-2x 2 . The set of possible values of...
- Let A and B be two events such that P( A B )= 16, P(A B)= 14 and P( A)= 14 where A = complementary of event A. Then A and B are
- Four flats are available in Krishna Apartment. Four persons visit the flats and each gives his choice to the promoter without...
- Three particles A, B and C start to move from three vertices of a triangle with the same uniform speed along the sides of the...
- A bag contains 12 tickets numbered 1,2,3, ,12. Five of them are drawn at random and arranged in the ascending order of their...
- Let E^c denote the complement of an event. Let E, F, G be pairwise independent events with P(G)>0 and P(E F G)=0. Then P(E^c F^c...
- STATEMENT-1: The probability of getting a tail most of the time in 10 tosses of a unbiased coin is 12 1 - 10! 2^ 10 ,5! ,5! ....
- STATEMENT-1: The probability of solving a new problem by 3 students are 12, 13 and 14 respectively. The probability that the...
Other chapters
Alcohols, Phenols and EthersAldehydes, Ketones and Carboxylic AcidsAlternating CurrentAminesAnatomy of Flowering PlantsAncient Indian HistoryAnimal KingdomApplication of DerivativesArea Under CurveAromatic Compounds and Aryl HalidesAtomsBasic Mathematics and LogarithmBinomial TheoremBiodiversity and ConservationBiological ClassificationBiomoleculesBiotechnology and Its ApplicationsBiotechnology: Principles and ProcessesBody Fluids and CirculationBreathing and Exchange of GasesCell Cycle and Cell DivisionCell: The Unit of LifeChemical Bonding and Molecular StructureChemical Coordination and Integration